# Evaluating World Models with LLM for Decision Making

### v1

@@ 2 Related Work
In this work, we provide a detail review of the related work, which is divided into three main categories, i.e., world models in decision making, LLMs as world simulators, world models in LLMs.
World Models in Decision Making.
@@
World models are actively explored by researchers to further improve the agent’s performance and the sample efficiency (Ha & Schmidhuber, 2018; Janner et al., 2019; Hafner et al., 2019; Schrittwieser et al., 2020). Dreamer (Hafner et al., 2019) is a practical model-based reinforcement learning algorithm that introduces the belief over states as a part of the input to the model-free DRL algorithm used. MuZero (Schrittwieser et al., 2020) is a remarkable success of model-based RL, which learns the world model and conduct the planning in the latent space. MuZero achieves the superior performances over other model-based and model-free RL methods. The world models trained in these methods are problem-specific and cannot be generalized to other problems, which motivates researchers to seek to more generalizable world models, e.g., LLMs as world models. The world model with LLM in (Xiang et al., 2023) is trained to gain the environment knowledge, while maintaining other capabilities of the LLMs. Dynalang (Lin et al., 2024) proposes the multimodal world model, which unifies the videos and texts for the future prediction in decision making.
LLMs as World Simulators.
@@
World simulators are developed to model the dynamics of the world (Bruce et al., 2024). LLMs serve as the world simulator due to their generalizability across tasks. Specifically, The LLMs (i.e., GPT-3.5 and GPT-4) is evaluated to predict the state transitions, the game progress and scores with the given object, action, and score rules, where these rules are demonstrated to be crucial to the world model predictions (Wang et al., 2024). The world models with LLMs in (Xie et al., 2024) need to additionally identify the valid actions. We move a step further to ask the world model to propose the potential actions to complete the tasks (Observation 2). Both methods mainly focus on the prediction of the state, which may be not suitable for the evaluation of the world model for decision making (Observation 1).
World Models in LLMs.
@@
The concept of world model also be explored in the deliberation reasoning of LLMs. Specifically, Reasoning via Planning (RAP) (Hao et al., 2023) leverage the planning methods (e.g., Monte Carlo Tree Search (MCTS)) with the world model with LLMs for plan generation and math reasoning, where LLMs need to predict the next state and the reward to guide the search. Tree of Thought (ToT) (Yao et al., 2023) implicitly leverage the LLMs as the world model to predict the next state and the reward for the search over different thoughts. Reason for future, act for now (RAFA) (Liu et al., 2023) combine the planning and reflection with the world model for complex reasoning tasks. However, these methods do not focus on the evaluation of the world models, and several interdependent modules are coupled with each other for completing the task (Observation 3).
3 Preliminaries
In this section, we will first introduce the preliminaries in decision making, the world model, search methods, and the recent practices that using LLM for decision making problems.
Markov Decision Process (MDP).
A decision making problem is usually represented as a Markov decision process (MDP) (Sutton & Barto, 2018), which is defined by the tuple $(S,A,T,R,\gamma)$, where $S$ is the state space, $A$ is the action space, $T:S\times A\rightarrow S$ is the transition dynamics, which specifies the next state $s^{\prime}$ given the current state $s$ and action $a$, $R:S\times A\rightarrow\mathbb{R}$ is the reward function, which specifies the agent’s reward given the current state $s$ and action $a$, and $\gamma$ is the discount factor. The agent’s policy is defined by $\pi_{\theta}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, parameterized by $\theta$, which takes the state $s$ as the input and outputs the action $a$ to be executed. The objective of the agent is to learn an optimal policy $\pi^{*}:=\arg\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}\right]$ is the expected return and $s_{0}$ is the initial state.
Large Language Models (LLMs).
Large Language models (LLMs) learn from text data using unsupervised learning. LLMs optimize the joint probabilities of variable-length symbol sequences as the product of conditional probabilities by $P(x)=\prod\nolimits^{n}_{i=1}P(s_{i}|s_{1},...,s_{i-1})$, where $(s_{1},s_{2},...,s_{n})$ is the variable-length sequence of symbols. LLMs with billions of parameters have demonstrated impressive capabilities in various NLP tasks (Touvron et al., 2023; OpenAI, 2023).
World Models.
The world model $\Omega$ is introduced to predict the dynamics of the environment, thus supporting the decision making process. Specifically, the world model is trained or prompted to predict the next state $s^{\prime}$, the reward $r$, and the terminal function $d$, given the current state $s$ and action $a$. The world model can be a neural network specially trained on the environments (Hafner et al., 2019; Schrittwieser et al., 2020), which cannot generalize across different environments. Therefore, world models with LLMs emerge as a promising methods due to the generalizbility of LLMs, where the prompting (Xie et al., 2024), in-context learning (Wang et al., 2024), and even fine-tuning methods (Xiang et al., 2023; Lin et al., 2024) are used to transform the LLMs to the world models.
4 World Models with LLMs for Decision Making
Figure 3: The next state prediction and reward/terminal prediction are considered in (Wang et al., 2024). With the two predictions, we can complete policy verification. We introduce the action proposal of the world model and the policy planning can be completed with the policy verification and action proposal.
In this section, we introduce the world model with LLM for decision making. Specifically, we will introduce the next state prediction, the reward and terminal prediction. Then, we will introduce how the world model will be used to complete the considered three main tasks, i.e., policy verification, action proposal, and policy planning. We provide the relationship between the three main tasks and the two kinds of predictions in Figure 3 for better understanding of the rationale behind the three tasks.
The world model considered in this work mainly follows the design in (Wang et al., 2024), where the representation of the state includes the objects in the environments and their properties. The prompts to the LLM, e.g., GPT-4o, also include the object rules, the action rules, and the score rules, which provides the necessary knowledge of the environments for the LLM to make accurate predictions. For the next state prediction, we ask the LLM to predict the state changes, i.e., the change of the objects’ properties, which is demonstrated to be efficient for the prediction (Wang et al., 2024). With the predicted state changes, we can recover the full state for further predictions. For the reward/terminal prediction, the LLM needs to predict three features: i) gameScore: the reward received from the environment, ii) gameOver: whether the task is terminated, and iii) gameWon: whether the task is successfully completed or not. For the rules used for the prediction, we refer to (Wang et al., 2024) for more details and the code for generating the prompts is also provided in Appendix B.1.
@@ 4 World Models with LLMs for Decision Making
Policy verification is one of the most straightforward way to apply the world model, which is motivated by Observation 1 to focus on the policy, rather than the predictions. The process for the policy verification is displayed in Algorithm 1. Specifically, given the environment env, the action sequence $\bm{a}$ with length $N$ to verify and the proportion of the action sequence to verify $\rho$, we will run the game for the first $\rho\cdot N$ steps (Line 1 of Algorithm 1), and leverage the world model to continue the last $(1-\rho)N$ steps (Line 1 of Algorithm 1). The returned results $r_{N},d_{N}$ will be compared with the true results from the environment to evaluate the performance of the world model for policy verification.
@@
The action proposal is a novel task for world model, based on Observation 2, which is not considered in previous work. Therefore, we design the prompt for the world model to generate top-$K$ potential actions to complete the tasks. Specifically, we follow the representation of the state in the next state prediction, with the additional information: i) the examples of actions, and ii) the previous actions. The code to LLM for the action proposal is displayed in Appendix B.2. One key issue for the action proposal is that the action generated by the world model may not be valid for the game at the current state. Therefore, given the predicted action $a^{\prime}$ and the set of possible actions to be executed at the current state $A^{\prime}$, we leverage the text-embedding model (OpenAI, 2022) to query the most similar actions with the cosine similarity, i.e., $a^{*}=\arg\max\{\texttt{emb}(a^{\prime},a),\forall a\in A^{\prime}\}$.
1 Given the env, the action sequence $\bm{a}$ to verify with $N=\texttt{len}(\bm{a})$, $\rho$ the proportion of $\bm{a}$ to verify, the world model $\Omega$
[…]
3 for $t\in\{1,2,...,N-1\}$ do
4       if $t<\rho\cdot N$ then $s_{t+1},r_{t},d_{t}=\texttt{env}(a_{t})$;
5       else $s_{t+1},r_{t},d_{t}=\Omega(s_{t},a_{t})$;
6
return $r_{N},d_{N}$
Algorithm 1 Policy Verification
1 Given the env, the action sequence $\bm{a}$ with $N=\texttt{len}(\bm{a})$, $\rho$ the proportion of $\bm{a}$ for planning, the world model $\Omega$, the planning sequence $\bm{a^{\prime}}=[]$
2for $t\in\{1,2,...,\rho\cdot N\}$ do
3       $s_{t+1},r_{t}=\texttt{env}(a_{t}),\bm{a}^{\prime}.\texttt{append}(a_{t})$
4
5for $t\in\{\rho\cdot N,\cdots,(2-\rho)N\}$ do
6       $a_{t}=\Omega(s_{t}),s_{t+1},r_{t},d_{t}=\Omega(s_{t},a_{t}),\bm{a}^{\prime}.\texttt{append}(a_{t})$;
7       if $d_{t}$ then break;
8
9
return $\bm{a}^{\prime}$
Algorithm 2 Policy Planning
@@
The policy planning task is motivated by Observation 3. The process of policy planning is displayed in Algorithm 2, which is built upon the policy verification and the action proposal (as depicted in Figure 3). Specifically, we execute the actions in the given $\bm{a}$ on the environment for $\rho N$ steps (Line 2 in Algorithm 2) and then plan for $2(1-\rho)N$ steps with the world model (Line 2 in Algorithm 2), where both the action to execute and the state transitions are generated by the world model. The returned action sequence $\bm{a}^{\prime}$ will be evaluated in the environment to verify the correctness. We note that only top-1 action is generated in Algorithm 2 for illustration, for the case where more actions are generated, we need to enumerate all possible outcomes or leverage advanced search methods, which will be tackled in future work.
5 Environments
Tasks.
We leverage the 31 diverse environments from (Wang et al., 2023) [^2] We note that there are 31 environments in (Wang et al., 2023), however, one environment, i.e., dish-washing, is used as the example for the world model, which is excluded for fair evaluation. with different tasks varying from daily tasks, e.g., washing clothes, to scientific tasks, e.g., forging keys, and different difficulty levels, i.e., steps to complete the tasks. The task suite is more diverse than the widely used environments, e.g., BabyAI (Chevalier-Boisvert et al., 2019) and MiniWob++ (Shi et al., 2017). A full list of the environments can be found in Appendix A.1.
Figure 4: Number of steps to complete the tasks
Rule-based Policies.
There are various randomness in the environments, including the specific tasks, e.g., the target color can be “orange”, “purple”, “green”, “black” in the paint task. However, for each task, only a single playthrough is provided in (Wang et al., 2024), which is not enough for a comprehensive evaluation of the world model for decision making. Therefore, we curate the rule-based policy for each environment and verify the correctness for 200 runs. The scripts for the rule-based policies are provided in Appendix A.2, which can help readers to understand the process to complete the tasks, as well as the complexities of tasks. We provide the statics of the number of steps to complete the tasks for 200 runs in Figure 4.333The names on the figure may differs from (Wang et al., 2023) for visualization, and please refer to Table 1 for correspondence.
6 Evaluations
In this section, we present the comprehensive evaluation of the world model for decision making over the diverse 31 environments on the three main tasks under various setttings. We use the advanced LLM models, i.e., GPT-4o and GPT-4o-mini, as the backbone LLM for the world model. We set the temperature of the LLM to be 0 to reduce the variance of the generation and all results are averaged over 30 runs.
6.1 Task I: Policy Verification
Evaluation Protocol.
Given the action sequence $\bm{a}$ generated by the rule-based policy, we leverage the world model to verify the last $\rho$ proportion of the policy, where $\rho\in\{0.25,0.5,0.75,1.0\}$. We note that when $\rho=1.0$, the world model will verify the full action sequence with only the initial observation of the environment. We say the verification of the policy is correct if all three features, i.e., gameScore, gameOver, and gameWon, are correct. We note that only correct policies are verified, as there are enormous wrong policies for an environment, which is useless for decision making. Furthermore, there would also be other action sequences to finish the tasks, where we cannot enumerate all policies to complete the tasks.
Evaluation Results.
The policy verification results are displayed in Figure 5. We observe that GPT-4o outperform GPT-4o-mini in most tasks and especially on the tasks which requires the domain knowledge, e.g., bandage, hang, and campfire. We also observe that with more steps of the verified policies, the performance gap between GPT-4o and GPT-4o-mini is increase. With larger proportion of the action sequences to verify, i.e., $\rho$ increase, the accuracy of the verification is decreased, which indicates that the accumulation of the errors in the world model, either on the next state prediction or the reward prediction, will influence the performance of the world model. This observation is consistent to the fact that the LLM may not perform well in long-term decision making tasks. We also observe that more steps to complete the tasks do not necessarily lead to the bad performance, which indicates that the domains of the tasks play a more important roles for the policy verification, i.e., for the tasks where the LLM has enough domain knowledge, the task would be easy even when the number of steps is large, e.g., conduct, stone, weigh and photo. We also provide the accuracy of the prediction of each feature in Appendix C, and we found that both GPT-4o and GPT-4o-mini performs wore for predicting the gameScore, while performs much better for predicting gameOver and gameWon, which indicates that the value prediction is more difficult for LLMs. During the experiments, we observe that both models usually output an empty dictionary of the verification result, which indicates the inabilities for the instruction following sometimes.
6.2 Task II: Action Proposal
Evaluation Protocol.
The action proposal requires the world model to generate the top-$K$ potential actions to complete the tasks, where $K\in\{1,2,3,5,10\}$. Specifically, given the action sequence $\bm{a}$ generated by the rule-based policies, we will let the world model to generate the potential actions with the states along with the path of $\bm{a}$ to complete the task. We say the action proposal is correct if the actions in $\bm{a}$ in the generated actions by the world model. The results of the accuracy are averaged over the steps over the action sequence and 30 runs of each environment. We also note that the action sequence $\bm{a}$ generated by the rule-based policy is not the only sequence to complete the task and we cannot enumerate all possible actions which can lead to the completion of the task. We note that the number of available actions in the environments is usually larger than 500, which brings difficulties to the traditional RL methods for training and indicate the necessity for the world model to generate the potential actions to facilitate the learning.
Evaluation Results.
The action proposal results are displayed in Figure 6. Overall, GPT-4o consistently outperforms GPT-4o-mini across different tasks and different values of $K$. With the number of steps to complete the tasks, GPT-4o maintains the better accuracy, while GPT-4o-mini shows a substantial drop of the accuracy. The performance gap between the two models is generally increased when the number of steps to complete the tasks increase. When $K=10$, the accuracy of the action proposal for GPT-4o is very high in most tasks, which indicates that the world model is capable to generate the relevant actions for completing the tasks while ignoring the irrelevant actions. Furthermore, we still observe that both models obtain lower values in the tasks requiring the domain knowledge, i.e., blood and conduct, which is consistent to the observation in (Wang et al., 2024) that LLMs (GPT-4) is more likely to make errors when scientific knowledge is needed. We also provide the step accuracy of the action proposal in Appendix D to illustrate the prediction of the relevant actions along with the steps. We observe that there are some key steps that has extremely low accuracies, which would be the most difficult steps for decision making.
6.3 Task III: Policy Planning
Evaluation Protocol.
The policy planning is based on the policy verification and the action proposal, as showed in Algorithm 2. Similar to the policy verification, we let $\rho\in\{0.25,0.5,0.75,1.0\}$ to vary the number of steps for the planning. We only consider the case with $K=1$, i.e., which generate the top-1 action with the given states predicted by the world model. Finally, we evaluate the planned policy $\bm{a}^{\prime}$ in the environment to verify the correctness. We note that when $K=1$, no advanced search method is needed, while when $K>1$, we cannot enumerate all possible outcomes for larger steps, e.g., 10. Besides, a critic is also needed to choose among the outcomes for verifying in the environments. Therefore, we only consider the case with $K=1$ and leave the case $K>1$ into future work.
Evaluation Results.
The policy planning results are displayed in Figure 7, where GPT-4o and GPT-4o-mini achieve comparable performance for the tasks with less steps and smaller values of $\rho$, e.g., $0.25$ and GPT-4o generally achieve better results in tasks with more steps. When the value of $\rho$ increases, the performance is generally decreasing. With the coupling of the policy verification and action proposal, we observe a more randomness of the performances of models over tasks and settings, which indicates that the necessity of decoupling the modules in decision making for comprehensive evaluation of the world model. During the experiments, we also observe the format errors of the outputs from both GPT-4o and GPT-4o-mini, which may interrupt the running of the experiments. Therefore, with the interaction of the different functionalities of the world model, the system is more unstable due to the unexpected in LLMs’ outputs.
Figure 7: The success rate of the world model to complete the tasks
7 Conclusions
World model is a key module for decision making and recent work leverage LLMs as the general world models. However, the evaluation of the world models with LLMs for decision making is far from satisfactory. In this work, we propose three main tasks, i.e., policy verification, action proposal, and policy planning, to evaluate the performance of the world model. We conduct the comprehensive evaluation of advanced LLMs, i.e., GPT-4o and GPT-4o-mini, on 31 diverse environments over the three main tasks under various settings. The key observations include: i) GPT-4o significantly outperforms GPT-4o-mini on the three main tasks, ii) the performance of the world model with LLM will be decreased for long-term decision-making tasks, and iii) the combination of different functionalities of the world model will brings unstability of the performance.
Appendix A Environments
A.1 Introduction of Tasks

### v2

@@ 6 Related Work
World Models in Decision Making. World models are actively explored by researchers to further improve the agent’s performance and the sample efficiency (Ha and Schmidhuber, 2018; Janner et al., 2019; Hafner et al., 2019; Schrittwieser et al., 2020). Dreamer (Hafner et al., 2019; 2025b; 2025a) is a practical model-based reinforcement learning algorithm that introduces the belief over states as a part of the input to the model-free DRL algorithm used. Trajectory Transformer (Janner et al., 2021) trains the transformer to prediction the next state and action as a sequence modeling problem for continuous robot control. MuZero (Schrittwieser et al., 2020) is a remarkable success of model-based RL, which learns the world model and conduct the planning in the latent space. MuZero achieves the superior performances over other model-based and model-free RL methods. The world models trained in these methods are problem-specific and cannot be generalized to other problems, which motivates researchers to seek to more generalizable world models, e.g., LLMs as world models. The world model with LLM in (Xiang et al., 2023) is trained to gain the environment knowledge, while maintaining other capabilities of the LLMs. Dynalang (Lin et al., 2024) proposes the multimodal world model to unify videos and texts for future predictions in decision making.
@@
LLMs as World Simulators. World simulators are developed to model the dynamics of the world (Bruce et al., 2024). LLMs serve as the world simulator due to their generalizability across tasks. Specifically, The LLMs (i.e., GPT-3.5 and GPT-4) is evaluated to predict the state transitions, the game progress and scores with the given object, action, and score rules, where these rules are demonstrated to be crucial to the world model predictions (Wang et al., 2024). The world models with LLMs in (Xie et al., 2024) need to additionally identify the valid actions. Recently, the Genie series (Bruce et al., 2024; Parker-Holder et al., 2024; Ball et al., 2025) learns controllable world models from unlabeled video data and enables users to interact with generated game-like environments through action inputs, followed by MatrixGame (Zhang et al., 2025; He et al., 2025) and GameCraft (Li et al., 2025). These work demonstrate the (multimodal) LLMs’ potential to serve as general world simulators. We move a step further to ask the world model to propose the potential actions to complete the tasks. Both methods focus on next state prediction, which may be not suitable for the evaluation of world models for decision making.
@@
LLM-based World Models for Decision Making. The concept of world model also be explored in the deliberation reasoning of LLMs. Specifically, Reasoning via Planning (RAP) (Hao et al., 2023) leverage the planning methods (e.g., Monte Carlo Tree Search (MCTS)) with the world model with LLMs for plan generation and math reasoning, where LLMs need to predict the next state and the reward to guide the search. Tree of Thought (ToT) (Yao et al., 2023) implicitly leverage the LLMs as the world model to predict the next state and the reward for the search over different thoughts. Reason for future, act for now (RAFA) (Liu et al., 2023) combine the planning and reflection with the world model for complex reasoning tasks. Recent work considers LLM-based world models for web agents (Chae et al., 2025; Gu et al., 2024), game agents (Hafner et al., 2025b), and even autonomous driving (Liao et al., 2025). However, these methods do not focus on the evaluation of world models, and several interdependent modules are coupled for completing the task.
7 Discussions
Are (LLM-based) World Models Necessary? Despite the success of MuZero (Schrittwieser et al., 2020) and Dreamer (Hafner et al., 2025a; b), world models remain relatively unpopular in the research community. The main barriers are the added complexity of training separate world models and their lack of generalizability. Although LLM-based world models offer improved generalization capabilities, most current LLM agents operate without them. This raises an important question: are (LLM-based) world models necessary? We argue that world models will become critical components of LLM agents for several compelling reasons. First, as agents are deployed in real-world scenarios where actions carry significant consequences, e.g., high-stake scenarios, world models enable essential counterfactual reasoning and outcome prediction prior to execution. This capability allows agents to simulate potential consequences without incurring the costs and risks of trial-and-error in physical environments. Second, world models enhance interpretability by explicitly modeling future states and outcomes, thereby making agent behavior more auditable and transparent to human stakeholders. Third, world models facilitate sample-efficient learning by enabling agents to plan and refine strategies through internal simulation rather than requiring extensive real-world interaction. Most critically, world models will play an essential role in safety assurance and governance, serving as a foundational safeguard in the development of safe artificial general intelligence (AGI). As we advance toward increasingly autonomous and complicated AI systems, the capacity to accurately model and reason about environmental dynamics and action consequences becomes not merely with better performance, but also necessary for responsible AI deployment in the real world.
Limitations. There are several limitations of the current framework. i) Our framework relies on rule-based policies as ground truth, which represents only one valid solution path per task. Real-world decision making often admits multiple valid strategies. This single-path assumption may penalize world models that propose alternative but potentially successful approaches, leading to underestimation of their true capabilities. ii) For policy planning, we only evaluate $K=1$ due to computational constraints. This limitation prevents us from assessing the world model to maintain consistency across multiple predicted trajectories, which is important for tree search methods. Extending to K>1 with tree search is important future work to better demonstrate full planning capabilities. To facilitate the understanding, we provide a comprehensive discussion about the role of $K$ in Appendix A.5. iii) While we argue that decoupling evaluation of world model functionalities is critical, this approach may miss important interactions between components. The performance degradation in policy planning suggests that integrated evaluation still provides valuable insights. iv) Due to the limited budget, we only conduct the experiments on two widely used models. Our claims should be interpreted as applying to the tested models, and validation across diverse model families (including open-source models like Qwen-72B, Llama-70B, and closed models like Claude and Gemini) remains important future work.
Conclusion. In this work, we argue that LLM-based world models can make decisions solely through the combination of policy verification and action proposal capabilities, but require rigorous, decoupled evaluation to understand their true strengths and limitations. Our proposed bottom-up evaluation framework, i.e., policy verification, action proposal, and policy planning tasks, enables systematic assessment across 31 diverse environments spanning daily and scientific domains. Key findings reveal that: i) GPT-4o significantly outperforms GPT-4o-mini, particularly on domain knowledge-intensive tasks, ii) performance depends predominantly on critical steps rather than total step count, and iii) integrating multiple world model functionalities introduces instability that obscures capability differences between models. While our framework has limitations—including reliance on single-path ground truth policies and restricted evaluation of multi-action planning—it establishes a foundation for systematic world model assessment that shifts focus from general simulation accuracy to decision-oriented performance. As LLM-based agents advance toward deployment in high-stakes real-world scenarios, our evaluation framework becomes essential for ensuring their safe and effective operation in complex decision-making tasks.
Acknowledgments
This work was partially supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Proposal ID: 23-SIS-SMU-037). The work described in this paper was also partially supported by a grant from the Innovation and Technology Commission of the Hong Kong Special Administrative Region, China (Project No. GHP/391/22).
Appendix A Frequently Asked Questions (FAQs)
A.1 Why World Models and Why World Models with LLMs?
i) Generalization to novel tasks: World models have demonstrated impressive transfer learning abilities (Byravan et al., 2020), allowing agents to adapt to previously unseen scenarios by leveraging their learned understanding of world dynamics. This generalization capacity is particularly valuable in robotics and control applications where agents must handle diverse situations (Robey et al., 2021; Young et al., 2023). ii) Efficient planning: The predictive capabilities of world models enable the sophisticated planning algorithms without any online interactions (Sekar et al., 2020; Hamrick et al., 2021; Schrittwieser et al., 2020). By simulating possible futures, agents can evaluate different action sequences and select optimal strategies without requiring actual interaction with the environment. This “imagination” or “mental simulation” capability dramatically improves sample efficiency and safety in decision-making. iii) Offline learning: World models have proven especially valuable in offline reinforcement learning settings (Schrittwieser et al., 2021; Yu et al., 2020; 2021), where agents must learn from pre-collected datasets without direct environment interaction. The ability to learn accurate dynamics models from historical data has opened new possibilities for training agents in scenarios where online interaction is impractical or costly.
A.2 More Justifications of the Proposed Tasks
In this section, we will provide a mode detailed justification of the three tasks proposed in this work.
@@ A.2 More Justifications of the Proposed Tasks
Policy verification can be viewed as a generalization of the next state prediction. Instead of focusing on the accuracy of the one-step prediction about the next states and the reward/terminal prediction, which is considered in most previous works, policy verification may accumulate the multi-step predictions and judging whether the given policy can complete the task or not. This task is more relevant to the world model for decision making, as if the world model can verify any given policy correctly, with the enough number of sampling of the policy, i.e., action sequences, we can complete the task in the end.
@@
As observing in the ToT (Yao et al., 2023), generating useful thoughts is critical which can significantly improve the performance. However, with multiple thoughts generated, we have to select one to executed. We can test these thoughts in the environments, however, this is not always doable. Therefore, building a world model is the straightforward way to do this. On the other hand, action proposal is necessary for the world model as the game engine to guide the fresh players to complete the game. With increasing the number of recommend actions, the difficulty of the action proposal is decreased. However, this task is not considered in the previous work for the world model. We believe that this task should be an important task to evaluate the world model for decision making.
@@
Policy planning is a combination of policy verification and action proposal. Conceptually, if the world model performs well on both tasks, we can obtain the policy with world model only and no actor is needed. This can help us to understand the world model through decoupling the world model with any other modules. Besides, this planning task is consistent with the System 2 thinking, i.e., with more time for the planning, the world model may find better policies.
A.3 The Objectives of This Paper
The primary objective of this paper is proposing the new evaluation tasks for the evaluation of the world models with LLMs for decision making.
  • Evaluating world model for decision making is difficult, given that the decision making tasks usually involves multiple steps of the predictions. Therefore, the one-step prediction tasks considered in most previous works is not suitable.
  • Instead of treating the world models as world simulators or supporting modules for the actors, we identify that the world model can solve the tasks solely with the combination of the policy verification and action proposal. Therefore, the world model should be researched with the same importance.
  • World models have traditionally been evaluated through a top-down approach, where complex systems are constructed to complete tasks, constraining analysis to high-level observations. By examining fundamental capabilities like policy verification and action proposal, we propose a bottom-up evaluation that enables more systematic and granular assessment of world models.
A.4 Selection of Backbone LLMs
  • We need to evaluate the three novel tasks over 31 environments, and each with 30 runs, we roughly use 3000 dollars for all the experiments for GPT-4o and GPT-4o-mini. Due to the limited budget, we cannot afford to test on Claude and Gemini.
  • For the open-sourced models, we test on some open-sourced models, e.g., Qwen 7B, and find that current open-sourced models still cannot generate the responses with correct formats, i.e., JSON. This brings difficulties for the evaluation.
A.5 The Role of $K$
The choice of $K$ in policy planning, i.e., the number of candidate actions proposed at each step, deserves careful discussion, as it governs the trade-off between world model isolation and practical planning performance, as well as the practical efficiency of applying world models for planning.
$K=1$ as a Principled Baseline.
Our main evaluation is with fixed $K=1$, which represents the most straightforward test of the world model: at each step, the model must commit to a single action with no opportunity for correction. This is a deliberate methodological choice rather than a computational limitation. When $K>1$, the planner must select among candidate sequences using a search strategy, e.g., beam search, majority voting, or MCTS, and success becomes a joint function of world model quality and search algorithm design. This reintroduces precisely the coupling problem identified in Observation 4.3, which our decoupled evaluation framework is designed to avoid. $K=1$ therefore provides the cleanest signal of the world model’s intrinsic planning capability, and establishes a meaningful lower bound.
What $K>1$ Would Look Like.
The action proposal results in Section 5.2 provide an indirect but informative window into $K>1$ planning behavior, since action proposal is essentially policy planning evaluated one step at a time. Two observations are relevant. First, at $K=10$, GPT-4o achieves near-perfect per-step coverage on the majority of environments, meaning the correct action is almost always present in the top-10 candidates. The $K=1$ planning bottleneck is therefore greedy commitment at critical steps, not a lack of world model knowledge. A simple $K>1$ selection strategy such as majority voting would be expected to substantially improve planning success rates, particularly on environments where a single bottleneck step has low $K=1$ accuracy but high $K=10$ coverage (e.g., make-campfire, space-walk). Second, for environments with long sequential dependencies (e.g., wash-clothes), increasing $K$ alone is insufficient: an error early in the sequence propagates regardless of how many candidates are generated later. In such cases, $K>1$ planning would need to be paired with policy verification as a subroutine, i.e., generating $K$ candidate continuations and verifying each before executing, which is a natural combination of our framework.
The Instability Concern.
Our experiment results in Subsection 5.3 show that combining world model functionalities already increases performance instability at $K=1$. $K>1$ planning would amplify this effect: each additional branch multiplies the surface area for compounding prediction errors and format failures. This suggests that the evaluation of $K>1$ planning should proceed carefully, with explicit ablations over search strategy and $K$ separately, rather than treating $K>1$ as a straightforward extension. We leave this systematic study to future work, but emphasize that our framework directly supports it, i.e., policy verification and action proposal are already evaluated as independent components, and their combination under $K>1$ search is a well-defined next step.
Summary.
$K=1$ is the starting point for understanding world model capability in isolation and $K>1$ is the next step for understanding world model capability in practice. The gap between them is bridged by search strategy design, which we treat as an orthogonal research question. Our results suggest that this gap is largely attributed to bottleneck steps rather than general knowledge deficiency, and that policy verification is the natural mechanism for exploiting $K>1$ candidates effectively.
A.6 Comparison with Decision Transformer and Trajectory Transformer
Decision transformer (DT) (Chen et al., 2021) trains the transformer to predict the action conditional on the experiences and the target reward or the goal. Language DT (LDT) (Gontier et al., 2023) extends DT to consider the text-based games and include the state prediction in the training as an auxiliary tasks. However, during the inference, i.e., decision making, the model still generates the action directly, which is not based on the world model because the state prediction is only used for training and not for acting.
Trajectory Transformer (TT) (Janner et al., 2021) also consider the decision making problem as a sequence modeling problem, where the transformer is trained to predict the state, the action and the reward. Compared with DT, TT is more related to the world model, where the state and reward prediction is used to generating actions and the search methods, e.g., beam search, is used. However, only the continuous robot control is considered in TT (Janner et al., 2021) and the trained TT model is domain and problem specific, which cannot generalize to other problems.
Recently, LLMs provide a promising way to build the general world model and the world model with LLMs emerge as a novel research field. However, most of these work focus on the single-step prediction and a comprehensive evaluate is needed. This work is inspired by TT and extends the insights from TT to the world model with LLMs for text-based games. Specifically, we consider the policy verification, the action proposal and the policy planning tasks, where the TT combines these tasks to generate the actions and only investigate the performance for the decision. Instead of only considering the performance of the decision makings, our three tasks provide a bottom-up analysis of the world model for decision making.
A.7 Broader Impact Statement
This work contributes an evaluation methodology for LLM-based world models towards AI safety and responsible development. Rigorous evaluation frameworks help prevent insecure deployment of vulnerable systems in high-stakes scenarios. However, we acknowledge potential dual-use concerns: improved world models could enhance autonomous agent capabilities, which could be misused for harmful automation. We emphasize that our work is evaluation-oriented and that substantial research on safety, robustness, and value alignment remains necessary before deploying world-model-based agents in real-world applications.
A.8 Code and Dataset Availability
The code and datasets can be access at: https://github.com/joannacyang/WorldModel_TMLR.
Appendix B Environments
B.1 Introduction of Tasks